OpenStudy (anonymous):

find second derivative of e^(8x) + 3cos(x)sin(x)

OpenStudy (anonymous):

\[e^(8x)+3\cos(x)\sin(x)\]

OpenStudy (anonymous):

so for the first derivative I get \[8e ^{8x}+3\cos^2x-3\sin^2x\] is that right so far?